A plane going to north with a speed of 150 m/s. It is blown by a strong wind at 500 m/s going to west what is the magnitude of plane's resultant velocity?
1
Expert's answer
2021-09-18T14:59:50-0400
"v= \sqrt{150^2+500^2}=522\frac{m}{s}"

"\theta=\arctan{\frac{150}{500}}=16.7\degree\text{ W of N}"